# Meta-learning of Gibbs states for many-body Hamiltonians with applications to Quantum Boltzmann Machines

### v1

@@ I Introduction
Generation of finite-temperature thermal states or Gibbs states of a given Hamiltonian is an integral task in state preparation using quantum computers. Gibbs states find applications in a variety of tasks such as quantum simulation [2], quantum machine learning [3, 4], quantum optimization [5], and the study of open quantum systems [6]. In particular, combinatorial optimization problems [5], semi-definite programming [7], quantum key distribution [8], quantum metrology [9], and training of Quantum Boltzmann Machines (QBMs) [3, 10, 11, 12, 13, 14, 15] are tackled by sampling from well-engineered Gibbs states.
Preparing arbitrary quantum states is generally a challenging task; in particular, finding the ground state of a Hamiltonian is known to be QMA-hard [16]. Preparing Gibbs states, especially at low and intermediate temperatures, is as hard as finding the ground state of that Hamiltonian [17]. Various methods propose preparing Gibbs states using classical and quantum computers [18, 6, 19, 20, 21, 22, 23, 24]. Some proposals that focus on quantum computers include quantum rejection sampling [25], quantum walk [26], and dynamics simulation [27, 28]. Although these methods may incur exponential cost in the worst case, they are efficient in special scenarios such as when the ratio between the partition functions of the infinite temperature state and the Gibbs state is at most polynomially large [6], and when the gap of the Markov chain describing the quantum walk is polynomially small [27, 26]. However, these algorithms suit fault-tolerant quantum computers as they require complex quantum subroutines like quantum phase estimation.
@@
Variational Quantum Algorithms (VQAs) that can be implemented efficiently on noisy intermediate-scale quantum (NISQ) devices, present an alternative to these methods. VQAs are a class of hybrid quantum-classical algorithms that optimize a loss function depending on the circuit parameters. The loss function is evaluated on quantum hardware, while parameter updates are computed on classical processors.
@@
Quantum Hamiltonian-Based Models and the Variational Quantum Thermalizer Algorithm, Verdon et al. [29] $4$ 0
Variational quantum Gibbs state preparation with a truncated Taylor series, Wang et al.[35] 11 1
Extending the Variational Quantum Eigensolver to Finite Temperatures, Selisko et al.[31] 4 0
Variational quantum simulation of the quantum critical regime, Shi et al. [40] 6 0
Training Quantum Boltzmann Machines with the $\beta$ Variational Quantum Eigensolver, Huijgen et al.[41] 10 0
Thermal state preparation of the SYK model using a variational quantum algorithm, Araz et al.[32] 12 0
Dynamic parameterized quantum circuits: expressive and barren-plateau free, Deshpande et al.[36] 10 0
Meta and Neural Network Meta-Variational Quantum Thermalizer, Our Work 8 8
@@ II.1 Variational Quantum Thermalizer (VQT)
Given a PQC represented by unitaries $U(\vec{\theta})$ with $m$ trainable angles, $\vec{\theta}\in\theta_{1},\theta_{2},...\theta_{m}$ acting on $\text{n}+\text{n}_{\text{ancilla}}$ qubits, the goal of VQT is to train the PQC to approximate the Gibbs state of this Hamiltonian as, $\rho(\vec{h})\approx\tilde{\rho}(\vec{\theta})=\text{Tr}_{\text{ancilla}}\left(\ket{\psi(\vec{\theta})}\bra{\psi(\vec{\theta})}\right)$, where |$\ket{\psi(\vec{\theta})}=U(\vec{\theta})\ket{\psi_{0}}$ is the state generated by the PQC from the initial state |$\ket{\psi_{0}}$. The exact loss function to be minimized using a classical computer to train this PQC is the Gibbs free energy given by,
@@
where $\langle H(\vec{h})\rangle=\text{Tr}[H(\vec{h})\tilde{\rho}(\vec{\theta})]$ is the expectation value of the Hamiltonian, $S=-\text{Tr}[\tilde{\rho}(\vec{\theta})\log\tilde{\rho}(\vec{\theta})]$ is the Von-Neumann entropy and $T=1/\beta$ is the temperature of the system. The minimization of $G(\vec{\theta},\vec{h})$ ensures that the system reaches the corresponding Gibbs state which is the state of minimal energy at finite-temperature. This is in analogy with the VQE where the expectation value of the Hamiltonian is minimized to find the ground state energy. In the limit $T\rightarrow 0$, Eq. (2) reduces to minimization of the expectation value of the Hamiltonian, which is the loss function in VQE.
However, unlike VQE, the mixed nature of the Gibbs state necessitates the use of ancilla qubits $(1~{}\leq~{}\text{n}_{\text{ancilla}}~{}\leq~{}\text{n})$ to prepare the approximate Gibbs state of given Hamiltonian. The required ancilla qubits can scale up further depending on how the loss function is evaluated on a quantum computer. For instance, Wang et al.[35] employs a truncated Taylor series expansion of order $3$ that needs a SWAP test to measure the approximate Von-Neumann entropy, thus requiring multiple copies of the quantum state and consequentially more ancillas. The schematic workflow of VQT is shown in Fig. 2(a).
@@ III.1 Finite-temperature collective optimization
In this work, we introduce the finite-temperature versions of Meta-VQE and NN Meta-VQE, known as Meta-VQT and NN Meta-VQT, which generates the Gibbs free energy profile, $G(\vec{h})$, and prepares the corresponding quantum Gibbs state, $\rho(\vec{h})$, of a parametrized Hamiltonian, $H(\vec{h})$. The workflow of our proposed Meta-VQT/NN Meta-VQT is shown in Fig. 2(b). The finite-temperature meta-algorithms are developed as generalized versions of the VQT algorithm explained in Sec. II.1. Here, the PQC, |$\ket{\psi(\vec{\theta},\vec{\phi},\vec{\omega})}$ (or neural network) is trained for multiple discrete input parameters ${\vec{h}_{train}}\in[\vec{h}_{1},\vec{h}_{2},...,\vec{h}_{n}]$ of the Hamiltonian, $H(\vec{h})$, so that for the optimal ansatz parameters, $(\vec{\theta}^{*},\vec{\phi}^{*},\vec{\omega}^{*})$ (or $\vec{\phi}_{nn}^{*})$, the quantum circuit is able to generate quantum Gibbs state for Hamiltonian parameters $\vec{h}_{test}\notin{\vec{h}_{train}}$. Suppose that $\vec{h}$ is a single-valued vector $h$ and $h_{train}\in(h_{1},h_{2},.....h_{n})$ is the training set with $h$ values arranged in ascending order. Then the meta-algorithms perform well for $h_{test}\notin h_{train}$ and $h_{1}\leq h_{test}\leq h_{n}$, whereas outside this range, its performance tends to decline. The number of distinctive trainable angles depends on the type of meta-algorithm, size of $\vec{h}$, the type of encoding and the class of PQC used for training.
@@
The loss function used is the sum of Gibbs free energy corresponding to each parameter in $\vec{h}_{train}$. For system sizes considered here, where the number of qubits ranges from $n=2~{}\text{to}~{}8$, the Gibbs free energy is evaluated on a classical computer after measuring the Gibbs state, $\rho(\vec{h})\approx\tilde{\rho}(\vec{\theta})=\text{Tr}_{\text{ancilla}}\left(\ket{\psi(\vec{\theta}^{*},\vec{\phi}^{*},\vec{\omega}^{*})}\bra{\psi(\vec{\theta}^{*},\vec{\phi}^{*},\vec{\omega}^{*})}\right)$, from the quantum circuit. For larger system sizes, where the size of the Gibbs state becomes too large to be stored in a classical computer, the Gibbs free energy itself is evaluated from the quantum circuits using a truncated Taylor series expansion of the entropy [35] or by applying the method described by Verdon et al. [29].
@@ III.1.1 Meta-Variational Quantum Thermalizer (Meta-VQT)
Figure 6: Gibbs state preparation of TFIM using Meta-VQT : The results of Meta-VQT algorithm run for different qubit sizes of TFIM is shown. The value of $\beta$ is chosen to be $1$. (a) shows the exact Gibbs free energy (dashed) evaluated in the interval $h\in[-2,2]$, and the one calculated after training from the variationally prepared Gibbs state for the parameters which are present (open circles) and absent (cross mark) in the training set. The relative error between exact Gibbs free energy and the Gibbs free energy evaluated using variationally prepared Gibbs state at $h_{test}$ with trained parameters of Meta-VQT is shown in (b). The relative error for all qubit sizes are low indicating the high efficiency of trained Meta-VQT in predicting the Gibbs free energy for test parameters outside the training set. This is also evident from (c) where the fidelity between exact Gibbs state and Gibbs state measured from the trained Meta-VQT for test parameters gives value between $0.93$ to $1$. The number of encoding and processing layers chosen for system sizes $n=2,3,4$ are $[2,2],[3,3]~{}\text{and}~{}[4,4]$ respectively.
@@ III.1.2 Neural Network Meta-Variational Quantum Thermalizer (NN Meta-VQT)
Figure 7: Gibbs state preparation of TFIM using NN Meta-VQT : The results of NN Meta-VQT algorithm for different qubit sizes of TFIM is shown. The value of $\beta$ is chosen to be $1$. (a) shows the exact Gibbs free energy (dashed) evaluated in the interval $h\in[-2,2]$, and the one calculated after training from the variationally prepared Gibbs state for the parameters which are present (open circles) and absent (cross mark) in the training set. The relative error between exact Gibbs free energy and the Gibbs free energy evaluated using variationally prepared Gibbs state at $h_{test}$ with trained parameters of NN Meta-VQT is shown in (b). The relative error for all qubit sizes are low indicating the high efficiency of trained neural network architecture in predicting the Gibbs free energy for test parameters outside the training set. This is also evident from (c) where the fidelity between exact Gibbs state and Gibbs state measured from trained NN Meta-VQT ansatz for test parameters gives value between $0.98$ to $1$ for $n=2,3$ and $4$. The number of SU2 hardware efficient and HVA ansatz layers in the processing part is chosen as $[2,2],~{}[3,3]$ and $[4,4]$ for system sizes $n=2,3$ and $4$ respectively.
Figure 8: Scaling of Meta-VQT and NN Meta-VQT:(a) shows the mean value with standard deviation of relative error between the Gibbs free energy evaluated for values of $h$ in $h_{test}$ from the trained meta-algorithms and the exact one for system sizes $2,4,6$ and $8$ qubits. The relative error increases for both Meta-VQT and NN Meta-VQT as system size increases to $8$, with NN Meta-VQT exhibiting lower value than Meta-VQT. (b) shows the mean value and standard deviation of fidelity evaluated between prepared and true Gibbs state at $h_{test}$ as a function of system size. The fidelity decreases for both Meta-VQT and NN Meta-VQT as system size increases to $8$, with a slighltly higher value for NN Meta-VQT than Meta-VQT for larger system sizes. For $n$ qubit Hamiltonian, the number of encoding (SU2 hardware efficient) and processing (HVA) layers considered for Meta-VQT (NN Meta-VQT) is $[n,n]$. The test set, $h_{test}$, over which the mean and standard deviation in relative error in Gibbs free energy and fidelity is evaluated are $40$ uniform points between $[-2,2]$.
@@ IV.1 Preparing Gibbs state of TFIM
In this section, Meta and NN Meta-VQT algorithms are deployed to generate the quantum Gibbs state of a parametrized transverse field ising model (TFIM), $H(h)=-J\sum_{i=1}^{N-1}\sigma_{i}^{x}\sigma_{i+1}^{x}-h\sum_{i=1}^{N}\sigma_{i}^{z}$, where the interaction strength $J=1$ and the Hamiltonian is parametrized as a function of transverse field, $h$. The inverse temperature $\beta$ is fixed as $1$, the number of ancillas is considered the same as the number of system qubits, and the initial learning rate is taken as $\eta=0.01$ for Meta-VQT and $0.001$ for NN Meta-VQT. The first set of experiments are run on three different system sizes $n=2,3~{}\text{and}~{}4$. The set of Hamiltonian training parameters $h_{train}$ considered are $10$ uniformly spaced points between $-2$ and $2$. The number of encoding and processing layers is chosen as $[2,2],~{}[3,3],~{}\text{and}~{}[4,4]$ for system sizes $n=2,3~{}\text{and}~{}4$, respectively. For all system sizes in NN Meta-VQT, the ansatz configuration has a feed forward neural network consisting of one input, one hidden and two output layers. The output layers from these neural networks provide angles for SU2 hardware efficient and HVA layers. For both Meta and NN Meta-VQT, the test set of Hamiltonian parameters, $h_{test}$ is considered as $40$ uniformly spaced points between $-2$ and $2$. This set also has points that are not present in the $h_{train}$.
Figure 6(a), shows the Gibbs free energy evaluated from the approximate Gibbs state measured from the trained Meta-VQT ansatz for different system sizes. The open circle corresponds to the variational Gibbs free energy evaluated at $h_{train}$ while the cross marker represents the one evaluated for $h_{test}$. The true Gibbs state evaluated analytically in the interval $[-2,2]$ is indicated as dashed lines. The relative error between true Gibbs free energy and the one evaluated from the approximate Gibbs state for the test parameter set is shown in Fig. 6(b). From Figs. 6 (a) and (b), we can conclude that the trained Meta-VQT can generate the Gibbs free energy profile with small relative error for all system sizes for the test set with just $10$ training points. Since the system size considered is relatively smaller, the quality of the prepared Gibbs state is further evaluated by analyzing its fidelity with the exact one as shown in Fig. 6(c). The fidelity stays in the range between $0.93$ to $1$, with a lower fidelity for larger system size.
@@
Figure 11: Preparation of Gibbs state in different regimes of Kitaev ring model : (a) shows the plots of Gibbs free energy evaluated from the approximate Gibbs state prepared using the trained Meta-VQT and VarQITE for different $h$ values of Kitaev ring Hamiltonian. For Meta-VQT the $h$ values considered are $40$ uniform points between $[0.7,1.2]$ including $20$ training points, while for VarQITE it is $20$ uniform points between $[0.7,1.2]$. The black dashed lines correspond to exact Gibbs free energy. The trained Meta-VQT generates Gibbs free energy profile more accurately than VarQITE even for $h$ values which are outside the training set. Additionally, (b) shows the fidelity calculated between the exact Gibbs state and the one approximated using both Meta-VQT and VarQITE. The fidelity for Meta-VQT remains close to one while that of VarQITE depends on the parameter values as well as the regimes of the Hamiltonian. The system size considered is $N=3$. The number of encoding and processing layers for Meta-VQT is $4$ SU2 hardware efficient and $1$ HVA respectively. The depth of SU2 hardware efficient layers in VarQITE is considered to be $5$.

### v2

@@ I Introduction
Preparation of quantum Gibbs states is of fundamental importance as they provide statistical description of quantum systems at finite temperature, which is the natural setting for almost all real-world physical systems. Beyond their foundational role in statistical mechanics such as the study of open quantum systems [2], Gibbs states have emerged as valuable resources in quantum technologies, underpinning applications in areas such as quantum simulation [3], quantum machine learning [4, 5, 6, 7, 8, 9, 10, 11], optimization [12], quantum metrology [13] and quantum cryptography [14]. Their broad utility, coupled with the intrinsic difficulty of preparing them efficiently [15], has motivated an extensive line of research aimed at developing algorithms and protocols for Gibbs state generation using classical and quantum computational techniques [16, 2, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26].
@@
While many of these existing quantum computing approaches are designed with fault-tolerant quantum hardware in mind, near-term devices are constrained by noise and limited circuit depth. Variational Quantum Algorithms (VQAs) that can be implemented efficiently on noisy intermediate-scale quantum (NISQ) devices, present an alternative to these methods. VQAs represent a class of hybrid algorithms that combine quantum and classical resources. The quantum device evaluates a loss function determined by the circuit parameters, and a classical optimizer updates these parameters iteratively.
@@
Quantum Hamiltonian-Based Models and the Variational Quantum Thermalizer Algorithm,
Verdon et al. [27]
44 0
Variational quantum Gibbs state preparation with a truncated Taylor series,
Wang et al.[33]
11 1
Extending the Variational Quantum Eigensolver to Finite Temperatures,
Selisko et al.[29]
4 0
Variational quantum simulation of the quantum critical regime,
Shi et al. [38]
6 0
Training Quantum Boltzmann Machines with the $\beta$ Variational Quantum Eigensolver,
Huijgen et al.[39]
10 0
Thermal state preparation of the SYK model using a variational quantum algorithm,
Araz et al.[30]
12 0
Dynamic parameterized quantum circuits: expressive and barren-plateau free,
Deshpande et al.[34]
10 0
Meta and Neural Network Meta-Variational Quantum Thermalizer,
Our Work
8 8
@@ II.1 Variational Quantum Thermalizer (VQT)
[…]
@@
where $\langle H(\vec{h})\rangle=\text{Tr}[H(\vec{h})\tilde{\rho}(\vec{\theta})]$ is the expectation value of the Hamiltonian, $S=-\text{Tr}[\tilde{\rho}(\vec{\theta})\log\tilde{\rho}(\vec{\theta})]$ is the Von-Neumann entropy and $T=1/\beta$ is the temperature of the system. The minimization of $G(\vec{\theta},\vec{h})$ ensures that the system reaches the corresponding Gibbs state which is the state of minimal energy at finite-temperature. The schematic workflow of VQT is shown in Fig. 2(a). This is in analogy with the VQE where the expectation value of the Hamiltonian is minimized to find the ground state energy. In the limit $T\rightarrow 0$, Eq. (2) reduces to minimization of the expectation value of the Hamiltonian, which is the loss function in VQE. However, unlike VQE, the mixed nature of the Gibbs state necessitates the use of ancilla qubits $(1~\leq~\text{n}_{\text{ancilla}}~\leq~\text{n})$ to prepare the approximate Gibbs state of given Hamiltonian.
II.1.1 Challenges in Gibbs Free Energy Evaluation and Scalable Approaches
Evaluating the exact Gibbs free energy becomes highly challenging for large system sizes. The main difficulty arises from the entropy contribution in Eq. 2, whose computation is generally intractable even on quantum computers for arbitrary large quantum states[45, 46]. To address this, several works have proposed algorithms that approximate the Gibbs free energy while maintaining scalability on near-term devices. Two notable approaches are:
  • The work by Wang et al.[33] expands the entropy in a Taylor series and truncates it at some finite order $K=2$ or $K=3$. The truncated series of entropy corresponds to evaluating higher-order overlaps between quantum states. Consequently, Gibbs free energy can be estimated as a linear combination of the system’s energy and these state overlaps, which are efficiently measurable on near-term quantum hardware. The trade-off is that evaluating higher-order state overlaps requires additional ancilla qubits.
  • Johannes Selisko et al.[29] introduced the quantum VQT (qVQT) algorithm, which employs two variational quantum circuits (VQCs) together with an intermediate measurement to prepare a mixed state that approximates the Gibbs state. The first VQC produces a probability distribution via intermediate measurement, enabling entropy estimation, while the second VQC evaluates the system’s energy. A classical optimizer iteratively tunes both parameter sets to minimize the Gibbs free energy. This hybrid quantum-classical scheme provides a resource-efficient and scalable route to Gibbs state preparation, adaptable to different ansatz families. Related approaches have also been applied by Verdon et al. [27] and J. Y. Araz [30] to prepare Gibbs states for models such as the 2D Heisenberg Hamiltonian and the Sachdev–Ye–Kitaev (SYK) model.
@@ III.1 Finite-temperature collective optimization
[…]
@@
[…]
@@ III.1.2 Neural Network Meta-Variational Quantum Thermalizer (NN Meta-VQT)
Figure 6: Gibbs state preparation of TFIM using Meta-VQT : The results of Meta-VQT algorithm run for different qubit sizes of TFIM is shown. The value of $\beta$ is chosen to be 11. (a) shows the exact Gibbs free energy (dashed) evaluated in the interval $h\in[-2,2]$, and the one calculated after training from the variationally prepared Gibbs state for the parameters which are present (open circles) and absent (cross mark) in the training set. The relative error between exact Gibbs free energy and the Gibbs free energy evaluated using variationally prepared Gibbs state at $h_{test}$ with trained parameters of Meta-VQT is shown in (b). The relative error for all qubit sizes are low indicating the high efficiency of trained Meta-VQT in predicting the Gibbs free energy for test parameters outside the training set. This is also evident from (c) where the fidelity between exact Gibbs state and Gibbs state measured from the trained Meta-VQT for test parameters gives value between $0.93$ to 11. The number of encoding and processing layers chosen for system sizes $n=2,3,4$ are $2,2],[3,3]~\text{and}~[4,4]$ respectively.
@@
Figure 7: Gibbs state preparation of TFIM using NN Meta-VQT : The results of NN Meta-VQT algorithm for different qubit sizes of TFIM is shown. The value of $\beta$ is chosen to be 11. (a) shows the exact Gibbs free energy (dashed) evaluated in the interval $h\in[-2,2]$, and the one calculated after training from the variationally prepared Gibbs state for the parameters which are present (open circles) and absent (cross mark) in the training set. The relative error between exact Gibbs free energy and the Gibbs free energy evaluated using variationally prepared Gibbs state at $h_{test}$ with trained parameters of NN Meta-VQT is shown in (b). The relative error for all qubit sizes are low indicating the high efficiency of trained neural network architecture in predicting the Gibbs free energy for test parameters outside the training set. This is also evident from (c) where the fidelity between exact Gibbs state and Gibbs state measured from trained NN Meta-VQT ansatz for test parameters gives value between $0.98$ to 11 for $n=2,3$ and $4$. The number of SU2 hardware efficient and HVA ansatz layers in the processing part is chosen as [2,2],[3,3][2,2],~[3,3] and $[4,4]$ for system sizes $n=2,3$ and 44 respectively.
Figure 8: Scaling of Meta-VQT and NN Meta-VQT:(a) shows the mean value with standard deviation of relative error between the Gibbs free energy evaluated for values of $h$ in $h_{test}$ from the trained meta-algorithms and the exact one for system sizes $2,4,6$ and $8$ qubits. The relative error increases for both Meta-VQT and NN Meta-VQT as system size increases to $8$, with NN Meta-VQT exhibiting lower value than Meta-VQT. (b) shows the mean value and standard deviation of fidelity evaluated between prepared and true Gibbs state at $h_{test}$ as a function of system size. The fidelity decreases for both Meta-VQT and NN Meta-VQT as system size increases to $8$, with a slighltly higher value for NN Meta-VQT than Meta-VQT for larger system sizes. For $n$ qubit Hamiltonian, the number of encoding (SU2 hardware efficient) and processing (HVA) layers considered for Meta-VQT (NN Meta-VQT) is $[n,n]$. The test set, $h_{test}$, over which the mean and standard deviation in relative error in Gibbs free energy and fidelity is evaluated are $40$ uniform points between [2,2][-2,2].
@@ IV.1 Preparing Gibbs state of TFIM
In this section, Meta and NN Meta-VQT algorithms are deployed to generate the quantum Gibbs state of a parametrized transverse field ising model (TFIM), $H(h)=-J\sum_{i=1}^{N-1}\sigma_{i}^{x}\sigma_{i+1}^{x}-h\sum_{i=1}^{N}\sigma_{i}^{z}$, where the interaction strength $J=1$ and the Hamiltonian is parametrized as a function of transverse field, $h$. The inverse temperature $\beta$ is fixed as $1$, the number of ancillas is considered the same as the number of system qubits, and the initial learning rate is taken as $\eta=0.01$ for Meta-VQT and $0.001$ for NN Meta-VQT. The first set of experiments are run on three different system sizes $n=2,3~\text{and}~4$. The set of Hamiltonian training parameters $h_{train}$ considered are $10$ uniformly spaced points between 2-2 and $2$. The number of encoding and processing layers is chosen as $2,2],~[3,3],~\text{and}~[4,4]$ for system sizes $n=2,3~\text{and}~4$, respectively. For all system sizes in NN Meta-VQT, the ansatz configuration has a feed forward neural network consisting of one input, one hidden and two output layers. The output layers from these neural networks provide angles for SU2 hardware efficient and HVA layers. For both Meta and NN Meta-VQT, the test set of Hamiltonian parameters, $h_{test}$ is considered as $40$ uniformly spaced points between 2-2 and $2$. This set also has points that are not present in the $h_{train}$.
Figure 6(a), shows the Gibbs free energy evaluated from the approximate Gibbs state measured from the trained Meta-VQT ansatz for different system sizes. The open circle corresponds to the variational Gibbs free energy evaluated at $h_{train}$ while the cross marker represents the one evaluated for $h_{test}$. The true Gibbs state evaluated analytically in the interval [2,2][-2,2] is indicated as dashed lines. The relative error between true Gibbs free energy and the one evaluated from the approximate Gibbs state for the test parameter set is shown in Fig. 6(b). From Figs. 6 (a) and (b), we can conclude that the trained Meta-VQT can generate the Gibbs free energy profile with small relative error for all system sizes for the test set with just $10$ training points. Since the system size considered is relatively smaller, the quality of the prepared Gibbs state is further evaluated by analyzing its fidelity with the exact one as shown in Fig. 6(c). The fidelity stays in the range between $0.93$ to $1$, with a lower fidelity for larger system size.
@@
Figure 11: Preparation of Gibbs state in different regimes of Kitaev ring model : (a) shows the plots of Gibbs free energy evaluated from the approximate Gibbs state prepared using the trained Meta-VQT and VarQITE for different $h$ values of Kitaev ring Hamiltonian. For Meta-VQT the $h$ values considered are $40$ uniform points between $[0.7,1.2]$ including 2020 training points, while for VarQITE it is 2020 uniform points between $[0.7,1.2]$. The black dashed lines correspond to exact Gibbs free energy. The trained Meta-VQT generates Gibbs free energy profile more accurately than VarQITE even for $h$ values which are outside the training set. Additionally, (b) shows the fidelity calculated between the exact Gibbs state and the one approximated using both Meta-VQT and VarQITE. The fidelity for Meta-VQT remains close to one while that of VarQITE depends on the parameter values as well as the regimes of the Hamiltonian. The system size considered is $N=3$. The number of encoding and processing layers for Meta-VQT is 44 SU2 hardware efficient and $1$ HVA respectively. The depth of SU2 hardware efficient layers in VarQITE is considered to be 55.